OpenStudy (anonymous):

I NEED HELP!!! WITH THIS QUESTION: Add and answer in simplest fractional form. 3/8 + -1/8

OpenStudy (anonymous):

\[ \frac{3}{8} -\frac{1}{8} = \frac{1}{4} \]